OpenStudy (anonymous):

Simplify 5 to the ^-2

OpenStudy (nincompoop):

what? no

OpenStudy (nincompoop):

\[ \sf \huge x^{-n} = \frac{1}{x^n} \]

OpenStudy (nincompoop):

\[ \sf example: \\ \huge 3^{-3} = \frac{1}{3^3}\\ \huge 6^{-2} = \frac{1}{6^2} \]